26. The demand for a particular part called SKU 005 is 1,500 units a year. The cost of one SKU 005 is $50.00. It costs $60.00 to place an order SKU 005, and the user of SKU 005 has a per year inventory carrying cost of 25% of unit cost. Assume that there are 250 working days in the year where SKU 005 is used.

a. What is the combined annual holding and ordering cost of an order size of 200 units for SKU 005?
Answer #1

Annual demand, D = 1500 units

Ordering cost, S = $ 60

Carrying cost, H = 50*25% = $ 12.5

a)

Given order size, Q = 200 units

Annual holding cost = H*Q/2 = 12.5*200/2 = $ 1,250

Annual ordering cost = S*D/Q = 60*1500/200 = $ 450

Combined annual holding and ordering cost = 1250+450 = $ 1,700